Nitromethane (CH-3NO_2) burns in air to produce significant amounts of heat. How much heat is produced by the complete reaction of 5.56 Kg of nitromethane?
1
Expert's answer
2019-11-21T06:26:37-0500

The molar mass of nitromethane,"m= (12+3(1) +14+2(16)) g\/mol" = 61g/mol

Thus, the number of moles n in 5.56 kg of nitromethane is:

"n =5560\/61" =91.15moles

Nitromethane (CH3NO2) burns in air to produce significant amounts of heat.

2CH3NO2(l)+3/2O2(g)-->2CO2(g)+3H2O(l)+N2(g) = -1418 kJ

According to the equation, 2 moles of nitromethane on combustion gives 1418 kJ of heat. Thus, the heat produced by the combustion of 91.15 moles of nitromethane:

Heat produced=1418kJ/2moles*91.15= 129250.7 J